Reports of 20434
MIF (Noise Budget)
takafumi.ushiba - 21:52 Friday 07 October 2022 (22415) Print this report
Transfer function measurement of BS TML to DARM

I measured the trasfer function from TML to DARM by using BS.
Figure 1 shows the result.
The data was stored at /users/Commissioning/data/NoiseBudget/TFs/2022/1007/TF_BS_TMLtoDARM_1007.xml

The slope is almost f^-2, which is consistent with theoretical one.
Noise projection to DARM from MICH feedback will be calculated soon.

Images attached to this report
CAL (Pcal general)
Shingo Fujii - 18:51 Friday 07 October 2022 (22414) Print this report
Periodic WSK calibration with GSK at U.Toyama

Date: 2022/10/7

With Dan Chen

We calibrated WSK (= IS between UToyama and KAGRA) with GSK calibrated by LIGO.

The measured ratio of WSK over GSK: 0.140267 +- 0.000101 [V/V]

Detailed report: link

MIF (ASC)
Terri Pearce - 17:55 Friday 07 October 2022 (22413) Print this report
DAC card, AI board and Converter board installed in ASC0 rack.
We installed AI board (S2011355) in row 15 of ASC0 rack and Dsub to BNC converter (S2214287) in row 27. AI is connected to the converter with 1 Dsub9 cable and 1 BNC is connected to the POP PZT1 Pitch controller. The AI is not connected to power (cable has been left near clean booth near ASC0 rack). We will check this connection works next week.

We installed a new DAC card (S2012039) in ASC0, slot 4. We tried DAC card S1605722 but we couldn't reboot with it so changed to this one and restarted successfully.
DGS (General)
satoru.ikeda - 17:30 Friday 07 October 2022 (22412) Print this report
Migration from k1script to k1script1

Overview
k1script is running on Debian8 OS.
We are planning to update it to Debian 10.
However, there are some resident softwares, so it is not possible to migrate.

Support
Migrated CRY and PEM cryocon daq and vacuum daq crontab.

Procedure
1. Updated Python from Python2 to Python3. Some errors were addressed.
 The old source code is kept as deb8.
2.Removed from crontab of k1script and added to crontab of k1script1.
 $ crontab -e
3.vacuum daq was newly created because key authentication was using it.

Next schedule
We will move the ones running from supervisorctl one by one.
We plan to ask you to take action as necessary.
 

Non-image files attached to this report
MIF (General)
takafumi.ushiba - 17:26 Friday 07 October 2022 (22411) Print this report
TEMP BIAS offset for PLLX was changed

Since PLL cannot be locked stably, I changed the temp bias offset for PLLX (K1:ALS-X_LASER_TEMP_BIAS_OFFSET) from -0.2 to -0.55.

VIS (BS)
lucia.trozzo - 17:06 Friday 07 October 2022 (22404) Print this report
Comment to IP geophone diagonalization. (22382)

Today I investigate the geophone residual coupling between DOF and I estimated the inter-calibration factors between the geophones and LVDTs along L, T, and Y directions.
Pic.1 shows the broadband residual coupling, as a function of the frequency, before and after changing the geophone align matrix.
 Looking at this plot, it is possible to note that the sensing couplings, especially the terms L2T, L2Y, T2L, and Y2T, are reduced.
The matrix seems to work fine.

Concerning the inter-calibration, a way to do that is to rescale the gain of the diagonalized geophone to match the amplitude of the TFs measured by the diagonalized LVDT_(L, T, Y). Keeping this idea in mind, I extrapolated the calibration factor of each geophone, at 0.2 Hz, calculating the ratio LVDT_(L, T, Y)/GEO_(L, T, Y).
Here are the values:
 

GEO_L= 1.0476
GEO_T= 1.1221
GEO_Y=1.2971

In pic. 2, Pic.3, and Pic.4, the comparison of the TFs before and after the calibration is shown. Looking at these plots, it is possible to see that the TFs between and GEOs are matching.

Note.
I haven't yet implemented these values in the MEDM, but I will as soon as the suspension is available.

Images attached to this comment
PEM (General)
tatsuki.washimi - 17:04 Friday 07 October 2022 (22410) Print this report
Comment to EX/EY Dsub15 -> MAG13, PTB110 (22353)

coherence with infrasound microphones

Images attached to this comment
IOO (OMC)
Michael Page - 16:56 Friday 07 October 2022 (22407) Print this report
Comment to Automation of OMC DCPD transimpedance gain filter - simulink/RCG modification (22357)

Ikeda, Michael

I checked to confirm the appropriate behaviour of the OMC_DCPD filter banks. When TIA is set to 100, FM8 should be on and FM9 should be off, and vice versa for TIA set to 400 (see figure 2). However, there was no change in the filter banks when I changed the TIA setting.

I tried replacing the switch in figure 3 with an AND gate (TIA_ZSW_OUT AND CONNECTED) but it didn't make a difference. Then I just removed the AND gate, sent CONNECTED to a terminator and put TIA_ZSW_OUT into B7 and B8, with the NOT gate still before B7. But that didn't make any difference either.

When I change the TIA setting in the OMC dock I can clearly see a change in the output signal of the DCPD, and using caget I can confirm that TIA_ZSW_OUT is going between "ON" and "OFF" (or 1 and 0 if using ezca read). So I guess the problem is somewhere between the model and the interface.

IOO (General)
satoru.ikeda - 16:54 Friday 07 October 2022 (22409) Print this report
Restarted models for k1omc, k1ascbeacon, k1sendbeacon, added Picimotor channel for WFS for beacon OMCASC

Performed the following tasks related to OMC

Request from michael-san

 Restarting the k1omc model was done with Mozumi.
 However, it did not work as expected.
 I will help again next time.

Request from Tanaka-san

 FE of k1ascbeacon and k1sendbeacon remained red.
 => I rebooted once and FE was restored.

 And I added a Picomotor channel used by WFS for beacon OMCASC.
 I will connect Picomotor next time.

Non-image files attached to this report
LAS (bKAGRA laser)
osamu.miyakawa - 16:39 Friday 07 October 2022 (22405) Print this report
Switched to HPL and PMC locked

I tried to switch the current KAGRA IR laser to the new high power laser (HPL). The new laser was injected in front of AOM.

Last week I moved the AOM 5cm to PMC direction and I moved 2.5cm more today to make a enough space for X-stage with a steering mirror.

The new bean has been  aligned to PMC and mode matching has been done with proper lenses and their positions.

Still alignment and mode matching are not perfect at all, but it was succeeded in locking PMC with the new HPL!

I needed to modulate to obtain error signal of PMC, I used the cable of current 15MHz signal and connected to a new 15MHz modulator in the HPL box.

After this switching test, I recovered the current KAGRA laser and PMC lock. Every thing should be back except for the position of AOM.

I will try a little more tuning for PMC and lock for IMC with HPL next week.

 

Images attached to this report
PEM (Center)
takaaki.yokozawa - 16:23 Friday 07 October 2022 (22408) Print this report
medm screen for new temperature monitors
After we installed the cryocon for monitoring the temperature of PSL and POP table and High Power Beam Damp(HPBD), we generated the very easy medm screen for monitoring them.
Images attached to this report
VIS (General)
satoru.ikeda - 16:23 Friday 07 October 2022 (22406) Print this report
Comment to Routine confirmation of DGS maintenance day (17831)

DGS Regular maintenance day(10/07)

All TestPoints were cleared before maintenance.

Stepper motor
 [Checked]
  Check if the power supply can be controlled by BIO
  Check if the script can be started
  Checked that the limit switches installed are green.
 [Check result]
  No problem
  BS is suspended for maintenance at other locations.

Release Sensor
 [Checked the Release Sensor.]
 Checked that Release Sensor is 0, 1: Seated.
 [Check result]
 ETMX, ETMY, ITMX, ITMY : No problem

Picomotor
 [Confirmation]
  PingTest is performed.
 [Result]
  Only known

 NG:
  BS(BF),SR2(IM)
   => Hung, rebooted and recovered.
  
  MCI,STM1,
  AS_WFS,POP_POM,POS_POM,POM1
  MCF(MCO,MCI)
   => Confirmation only, no recovery operation was performed.

 OK:
  ETMX,ETMY,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF),SR3(IM,BF,STM)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCE,IMMT1,IMMT2,OMMT1,OMMT2,OSTM
  REFL_WFS,POP,POP2,POS,POS2
  PCAL(EX1,EX2,EY1,EY2):Not checked because it is ON only when used

Temperature controller
 [Checked]
 Is the temperature acquisition working properly?

 [Check result] 
  Is the temperature acquisition working properly?
  => No problem

Precision air Processer
 [Check Contents]
 Is the temperature acquisition working properly?

 [Result]
  No problem

FunctionGenerator
 (Confirmation)
 Is there a ping response?
 Are all EPICS channels alive?
  [sitemap]-[Commissioning]-[ALS FG]-[Xarm],[Yarm].
 [Check result]
 X and Y PLLs are working properly.

HWP control PC
 (Confirmation)
 Check if it is possible to login to the control PC.
 (Result) 
 Both k1hwp0 and 1 are OK.
 

Images attached to this comment
VIS (OMM)
ryutaro.takahashi - 13:34 Friday 07 October 2022 (22403) Print this report
Check of OSTM suspension

I checked the OSTM suspension. I found the IM is touching the magnets in X+Y-.

Images attached to this report
MIF (ITF Control)
takaaki.yokozawa - 12:14 Friday 07 October 2022 (22402) Print this report
Preparation of the QPDs for OMMT2 trans
I prepared two QPDs for the OMMT2 trans monitoring.
Now, two QPDs are placed on the table of the -x side of the OMMT2 chamber.
Images attached to this report
VAC (General)
shoichi.oshino - 11:38 Friday 07 October 2022 (22401) Print this report
Preparation of PC for Q-mass measurement
I prepared PCs for Q-mass measurement.
For IX, I installed Qulee QCS of version 3.
For BS and EX, I installed Qulee QCS of version4.
IOO (OMC)
Michael Page - 10:02 Friday 07 October 2022 (22399) Print this report
OMC medm dock update

I updated some of the monitoring channels in the OMC dock ndscope popups. Specifically I changed the OMC LSC and ASC ndscope monitors to use _DQ channels. In the OMC Lock module, I also changed the TMS IR monitors from AOS-TMSX(Y)_IR_PDA1_OUTPUT to TMS-X(Y)_IR_PDA1_OUT_DQ.

Images attached to this report
DGS (General)
shoichi.oshino - 9:32 Friday 07 October 2022 (22400) Print this report
Package update
I updated several packages on k1detfs0.
VAC (SRM)
nobuhiro.kimura - 5:39 Friday 07 October 2022 (22398) Print this report
Result of vacuum leak test on SRM (Continued)
When: 06/Oct./2022
Who: Kimura, Yasui, M. Takahashi, Uchiyama and Miyoki
Task:
We performed a vacuum leak test on the SRM.
There were no leakage >1 x 10^-11 Pam^3/s except for flanges #3 and #3-4.
Please confirm position of the flanges #3 and #3-4 with attached drawing.
The flanges #3 and #3-4 are difficult to access and require attention to Oplev.
Need to discuss with the the involved personnel to inspect flanges #3 and #3-4 for leaks.
Images attached to this report
CAL (General)
takahiro.yamamoto - 2:54 Friday 07 October 2022 (22396) Print this report
Comment to DARM calibration was updated (22366)
I found one or two mistakes in the polarity of the calibration factors.

One is just my mistake to treat DARM2 gain as 1.0 in DARM model. It is -1.0 in correctly.

Another one is about relation between +L of DARM and +L of VIS.
I will check it more carefully tomorrow.

IOO (OMC)
yohei.nishino - 0:19 Friday 07 October 2022 (22394) Print this report
Transimissivity estimation

We estimated the transmissivity of the OMC.

Since we can not scan the cavity length, we try to lock every transmitted mode and check its power. This measurement is done with 100 Ohm transimpedance.

Incident power ~ 81 [mW]

Transmission ~ 1 + 2 + 5 (TEM02 of carrier) + 0.4 + 0.2 (TEM00 of RF sideband) + 15 (TEM10) + 50 (TEM00 of carrier) = 73.6 [mW]

total power ratio ~ 73.6/81 = 90.8 %

HOMs with small power were hard to identify the mode.

The remaining 10 % of light could be lost by mirror contamination. For a more precise estimation, we need to scan the cavity in a more stable situation, but it couldn't be realized without drawing vacuum.

DetChar (General)
hirotaka.yuzurihara - 21:49 Thursday 06 October 2022 (22393) Print this report
Script to show the improvement of KAGRA's sensitivity during O4 commissioning.

I prepared the script (jupyter notebook) to show the improvement of KAGRA's sensitivity during O4 commissioning.

The plots (by using sensitivity on Sep. 22) and jupyter notebook were uploaded on JGW doc. Please use it after editing as you want. Sorry for late post.

MIF (ASC)
hirose.chiaki - 21:45 Thursday 06 October 2022 (22391) Print this report
Comment to ASC Model POP WFS Update (22362)

[Terri Chiaki]

We updated MEDM screen(ASC_OVEWVIEW.adl) to support POP_PZT for DC centering.
detail : 

  • In the actuator corners, we created "POP_PZT{1,2}" filter bank in both PIT and YAW directions.
    (The ON/OFF button should also work.)
  • We added 2 line of matrix in both PIT and YAW directions. 
    For this reason, the matrix of basic was created on this file.(./opt/rtcds/userapps/release/isc/common/medm/fmX/ASCMTRX{P,Y}_Kissel.adl)
    We set the address to the matrix using emacs so that pressing this button would display the screen we wanted.
Images attached to this comment
MIF (General)
takafumi.ushiba - 21:20 Thursday 06 October 2022 (22392) Print this report
Comment to FPMI was NOT locked after alignment recovery (22389)

Strictly speaking, we can lock FPMI with 3F but cannot hadover 3F signals to 1F signal due to the beam lost at POP PDA1, which is used for MICH lock.
Since CARM and DARM signals can be obtained at REFL and AS, in-vac alignment should be fine, so the suspicious point is POP table. 

MIF (ASC)
Terri Pearce - 20:29 Thursday 06 October 2022 (22390) Print this report
POP WFS alignment and RFQPD no signal
[Keiko, Chiaki, Terri]

POP WFS PZT1 and PZT2 have been connected to the controller via BNC cables and labelled.

POP WFS optics have been aligned with the ITMX direct reflection to show an ndscope signal around 200 counts (with around 2W laser power) in the SPOL PD (after RBS6), channel name K1:LSC-POPF_SPOL_DC_OUT16, as well as a beam in the SPOL CCD camera. We could not align to the PPOL PD (K1:LSC-POPF_PPOL_DC_OUT16) or PPOL CCD as the beam is very weak here, following the TFP, and was difficult to see.

There was no DC output from QPD1 and 2. We tried a spare QPD with the same Dsub15 cable and there was also no response and then we tried another cable and there was still no response. We are unsure why, as we couldn't find a problem in the signal chain.
Images attached to this report
MIF (General)
masahide.tamaki - 20:05 Thursday 06 October 2022 (22389) Print this report
FPMI was NOT locked after alignment recovery

After alignment work, we tried to lock FPMI, but we couldn't lock because there was no light in the POP PDA1.
This should be checked tomorrow.

► Trasmitted light powers were as below.

  • GRX ~ 1.41
  • IRX ~ 8.5
  • GRY ~ 0.53
  • IRY ~ 38.5
Images attached to this report
Comments to this report:
takafumi.ushiba - 21:20 Thursday 06 October 2022 (22392) Print this report

Strictly speaking, we can lock FPMI with 3F but cannot hadover 3F signals to 1F signal due to the beam lost at POP PDA1, which is used for MICH lock.
Since CARM and DARM signals can be obtained at REFL and AS, in-vac alignment should be fine, so the suspicious point is POP table. 

Search Help
×

Warning

×